If the perimeter of a rectangle is 7 3/5 units. What is the value of x? The rectangle has 2 sides as x and 2 sides as x+4/5
Answer by addingup(3677)

The rectangle has two short sides the problem calls x, these are the width
The rectangle has two long sides we will call L for length
And the problem says the long sides are 4/5 longer than x
We have:
2x+2L = 7 3/5 and since L= x+4/5:
2x+2(x+4/5)= 7 3/5 Multiply on the left
2x+2x+8/5= 7 3/5 Convert this mixed number to fraction
2x+2x+8/5= 38/5 Subtract 8/5 on both sides and add x on left
4x= 30/5 Simplify, 30 goes into 5 6 times (5*6= 30)
4x= 6 Divide both sides by 4
x= 6/4 Four goes into 6 one time and you have a remainder of 2
x= 1 2/4 2 Two and 4 are both divisible by 2:
x= 1 1/2